alones (alones):

How do i properly locate that :3

alones (alones):

3a^2 · 2a^3

OpenStudy (anonymous):

idk

alones (alones):

umm okay :3

OpenStudy (anonymous):

you're welcome, I'm expecting about a hundred medals

OpenStudy (anitas.):

what do you mean by locate?

OpenStudy (jdoe0001):

locate? what the?

OpenStudy (anitas.):

^

alones (alones):

mhm what the..finnish it xP Place the indicated product in the proper location on the grid.

alones (alones):

@Directrix help me :3 i'm stuck

Directrix (directrix):

Is this the question? Place the indicated product in the proper location on the grid. 3a^2 · 2a^3

alones (alones):

Oh yea..s orry if it took your time to find it

Directrix (directrix):

3a^2 · 2a^3 = 3*2 * a^2 * a^3 =

Directrix (directrix):

What is 3 times 2 =

Directrix (directrix):

@AloneS

alones (alones):

Oh sorry... 3*2 is 6 ._>

Directrix (directrix):

So, 3*2 * a^2 * a^3 = 6 * a^2 * a^3 If you multiply, a*a * a * a *a, how many factors of a are you multiplying?

Directrix (directrix):

@AloneS

Directrix (directrix):

Count them: a*a * a * a *a,

alones (alones):

it's 5

Directrix (directrix):

a*a * a * a *a = a^5 which is what a^2 * a^3 equals.

Directrix (directrix):

3*2 * a^2 * a^3 = 6 * a^2 * a^3 = 6 * a^5

alones (alones):

so a^5?

Directrix (directrix):

I don't see a grid.

alones (alones):

Oh i don't either :3

Directrix (directrix):

This is the answer to the posted problem: 6 * a^5 I don't know what you mean by grid because I cannot see your computer screen.

Directrix (directrix):

Look at what you posted above: Place the indicated product in the proper location on the grid.

alones (alones):

LOL xP I wish ummm mayeb the graph eh let me screenshot that

Directrix (directrix):

Probably the answer grid on the answer section of the computer program.

alones (alones):

okay so here

alones (alones):

DAANG!!That’s the wrong one -.- here sorry @Directrix it's laging for me

Directrix (directrix):

That is how I would enter the answer in the grid.

alones (alones):

Oh okay thanks...so i think i'll get this right

Directrix (directrix):

Enter the product ---> That is the enter answer as you have entered it.

alones (alones):

Oh i was breaking my head UGH i though i had to write whole th equation there -.-

Directrix (directrix):

This thread is unlucky. Close it and move on, I say.

alones (alones):

Oh yea terriable one thanks though<3